025 M Sr(NOh M Sr(NO,) 3)2- Calculate the molar solubility of Sr,(PO)h in 0.025 M S 2. -3 mol/L at a certain temperature. 19. The molar solubility of Zn(OH)2 is 5.7 x 10 Calculate the value of Ksp for Zn(OH)2 at this temperature
Answer #1

18. In order to answer this question, we need to know the molar solubility of Sr3(PO4)2 which is 1.0*10-31. We need to write down the dissociation equation and set up the ICE chart.

Sr3(PO4)2 ç=====è 3 Sr2+ + 2 PO4

Let x mole/L be the molar solubility of Sr3(PO4)2. Then the solubility of Sr2+ and PO43- will be 3x and 2x as per the dissociation equation. Also, we have 0.025 M Sr(NO3)2 which furnishes 0.025 mole/L Sr2+. Therefore,

Ksp = [Sr2+]3[PO43-]2 = (0.025 + 3x)3(2x)2 where Ksp is the solubility product of Sr3(PO4)2. We need to make an assumption here: x << 0.025 and hence [Sr2+] = 0.025. Therefore,

1.0*10-31 = (0.025)3(2x)2 = (1.5625*10-5).(4x2) = 6.25*10-5.x2

=è x2 = 1.6*1027

=è x = 4.0*10-14

The molar solubility of Sr3(PO4)2 in 0.025 M Sr(NO3)2 is 4.0*10-14 M (ans).

19. The molar solubility of Zn(OH)2 is 5.7*10-3 mol/L at a certain temperature. The dissociation of Zn(OH)2 occurs as

Zn(OH)2 ç==è Zn2+ + 2 OH-

[Zn2+] = 5.7*10-3 mol/L

[OH-] = 2*5.7*10-3 mol/L; this is as per the dissociation equation. Therefore,

Ksp = [Zn2+][OH-]2 = (5.7*10-3).(2*5.7*10-3)2 = 7.407*10-7 ≈ 7.41*10-7 (ans).

The molar solubility of CuCl in water is 1.0*10-3 mol/L. The dissociation of CuCl takes place as

CuCl ç===è Cu+ + Cl-

[Cu+] = 1.0*10-3 mol/L

[Cl-] = 1.0*10-3 mol/L; this is as per the dissociation equation. Therefore,

Ksp = [Cu+][Cl-] = (1.0*10-3).(1.0*10-3) = 1.0*10-6 (ans).